OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B is a square pyramid with a base side length of 36 inches and a height of 24 inches. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (anonymous):

please somone

OpenStudy (anonymous):

ok so the volume for a square pyramid is Volume = (base side length) (height/3)

OpenStudy (anonymous):

for pyramid A we would get a volume of 384 cubic inches. for pyramid B we would get a volume of 10368 cubic inches

OpenStudy (anonymous):

I meant (base side length)^2

OpenStudy (anonymous):

so \[(12)^2(8/3)=384\] \[(36)^2(24/3)=10368\]

OpenStudy (anonymous):

10368/384=27